# Enhancing Large Language Models for Time-Series Forecasting via Vector-Injected In-Context Learning

### v2

@@ C.1. Variable look-back length
We systematically evaluate our method under multiple look-back length configurations to examine its sensitivity and robustness to changes in the historical window. In all experiments, the prediction length is set to 96. The results are shown in Figure 8: although the metrics fluctuate across settings, the overall trend is clear— as the look-back length increases, the model leverages richer historical context and achieves steady improvements in predictive accuracy. However, excessively long windows slow inference, and the performance gains taper off once the look-back length exceeds 672. Balancing predictive benefits against computational cost, we set the look-back length to 672.
@@ C.2. Number of Examples Used
In this section, we evaluate how using different proportions of the training data (0%–20%) as examples affects the model’s forecasting performance. The experimental setup fixes the look-back window at 672 and the forecast horizon at 96; results are shown in Fig. 9. We observe that 1) performance improves as the proportion of training data used for examples increases, and 2) once the proportion exceeds 5%, the growth in gains slows noticeably. Accordingly, we use 5% of the training data as examples to construct the context vector, striking a balance between performance and efficiency. Note that all examples are compressed into a single context vector, so the number of examples does not affect the model’s inference speed.
@@ Appendix D Inference Efficiency Comparison of Vector-Injected and Traditional ICL
Appendix E Additional Related Work
E.1. Residual Stream
Recent academic applications of the residual stream have primarily focused on NLP and multimodal settings. (Merullo et al., ) find that vectors constructed from positive/negative texts or learned via gradient descent, when injected into the residual stream at inference time, can stably steer an LLM’s content and style. (Hua et al., 2025) construct activation data on the residual stream for objects labeled “factual/hallucinatory,” train sparse autoencoders (SAEs) to extract specific features, and then inject them into the residual stream, effectively reducing hallucinations in multimodal models. (Allbert et al., 2024) constructs and injects directions/vectors related to personality traits in the residual stream, enabling controllable enhancement and composition of LLM personality attributes. (Govindan et al., 2025) propose injecting vectors into the residual stream at inference time to impart temporary memory to LLMs and thus influence their response style. Inspired by these methods, we propose injecting the context vector generated from time-series exemplars into the residual stream to intervene in the LLM, making it better suited for time-series forecasting (TSF) tasks.
E.2. ICL for large models for TSF
Several studies (Lu et al., 2024; Das et al., 2024) apply in-context learning (ICL) to large models for time-series forecasting (TSF). These large models are pretrained on datasets comprising hundreds of millions of time series, and their primary concern is how such large models for TSF transfer to the downstream TSF task. Moreover, these approaches typically realize ICL by explicitly concatenating the examples into the prompt, which results in overly long prompts. In contrast, this paper focuses on transferring large language models (LLMs) to TSF tasks and adopts a vector-injection ICL to avoid excessively long prompts. Overall, although both lines of work employ ICL, they differ substantially in their issue focus and concrete implementations.

### v3

@@ C.1. Variable look-back length
We systematically evaluate our method under multiple look-back length configurations to examine its sensitivity and robustness to changes in the historical window. In all experiments, the prediction length is set to 96. The results are shown in Figure 8a: although the metrics fluctuate across settings, the overall trend is clear— as the look-back length increases, the model leverages richer historical context and achieves steady improvements in predictive accuracy. However, excessively long windows slow inference, and the performance gains taper off once the look-back length exceeds 672. Balancing predictive benefits against computational cost, we set the look-back length to 672.
@@ C.2. Number of Examples Used
In this section, we evaluate how using different proportions of the training data (0%–20%) as examples affects the model’s forecasting performance. The experimental setup fixes the look-back window at 672 and the forecast horizon at 96; results are shown in Fig. 8b. We observe that 1) performance improves as the proportion of training data used for examples increases, and 2) once the proportion exceeds 5%, the growth in gains slows noticeably. Accordingly, we use 5% of the training data as examples to construct the context vector, striking a balance between performance and efficiency. Note that all examples are compressed into a single context vector, so the number of examples does not affect the model’s inference speed.
@@ Appendix D Inference Efficiency Comparison of Vector-Injected and Traditional ICL
Appendix E More Method Analysis
E.1. Effectiveness of Aggregation
In LVICL, we aggregate feature vectors from multiple samples into a single context vector to capture features shared across samples. We hypothesize that such shared features are more relevant to the prediction task than single-sample representations. To validate this, we measure the average mutual information (MI) between the aggregated context vector constructed from varying numbers of samples from 1 to 5000 and the predicted series. There is no overlap between the predicted series and the samples participating in the aggregation. MI quantifies statistical dependence, with higher values indicating stronger correlation. The predicted series (length $P$) is transformed via a trained embedding layer to match the context vector’s dimensions, after which their MI is computed using the histogram method. Further details on the histogram method are in (Steuer et al., 2002). The results are shown in Figure 10a, where each point averages 20 random trials. The results show that both mean and adaptive weighted aggregation yield significantly higher MI than single-sample representations, and the MI increases with the number of aggregated samples. These results indicate that cross-sample shared features are more strongly correlated with the prediction task.
E.2. Comparison of Aggregation Methods
We compare the final performance of the model using mean aggregation in Eq. 7 and adaptive weighted aggregation. The experimental results are shown in Figure 10b. The results indicate that mean aggregation outperforms adaptive weighted aggregation.
E.3. Effectiveness of Context Vector Adapter
We compute the MI between the context vector and the predicted series, both before and after processing by the Context Vector Adapter. The results show that the values are comparable, remaining at approximately 0.56. However, the ablation study in Figure 4 demonstrates that using the context vector processed by the adapter yields a significantly lower prediction MSE. This indicates that the adapter effectively filters out prediction-irrelevant redundant information from the context vector while preserving key prediction-related features, thereby enhancing the model’s overall performance.
Figure 10. (a) Average MI between the aggregated context vector and the predicted series under different numbers of aggregated samples. (b) Comparison of two aggregation methods.